OpenStudy (anonymous):

Convert the following common fractions to decimals. Round to three decimal places. You may use a calculator. 1/3 =

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }=0.333\]

OpenStudy (anonymous):

..0303030303030303030 on and on and on

OpenStudy (anonymous):

oh i mean .333333333333333333333333 repeating so just write it as 

OpenStudy (anonymous):

@ hwtrain please explain to me how u got that answer.

OpenStudy (anonymous):

use a calculator

OpenStudy (anonymous):

and then what?

OpenStudy (anonymous):

well do you know what a fraction is?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

as in what is a fraction?

OpenStudy (anonymous):

ok so what is it?

OpenStudy (anonymous):

1/2 is a fraction

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }=\frac{ x }{ 100 }\]cross multiply to get 100 = 3x Divide by 3 to get x by itself 33.33333333 = x Now just convert it to only Decimal form and round 0.333

OpenStudy (anonymous):

a fraction is a division so you just use in a calculator and write the numerator which is one in this case and then division sign then the denominator which is 3 so then you should get .333